5(((x + 1 = 3)))
I hope that I typed this correctly. 5 then square root sign then x plus 1 equals 3.

Answer by stanbon(75887)

5(((x + 1 = 3)))
I hope that I typed this correctly. 5 then square root sign then x plus 1 equals 3.
---------------
5sqrt(x+1) = 3
sqrt(x+1) = 3/5
x+1 = (3/5)^2
x = 9/25 - 25/25
x = -16/25
======================
Cheers,
Stan H.